# DVMNet: Computing Relative Pose for Unseen Objects Beyond Hypotheses

### v1

@@ 4.3 Relative Object Pose Estimation on GROP
Recently, a new benchmark called GROP for relative pose estimation of unseen objects was introduced in [52]. This benchmark comprises two datasets, i.e., Objaverse [9] and LINEMOD [15]. Both synthetic and real images with diverse object poses are considered. We develop experiments on these two datasets, following the same setup as described in [52]. More concretely, the synthetic images are generated by rendering the object models of the Objaverse dataset from different viewpoints [20]. Several sequences of calibrated real images that depict 13 texture-less household objects are provided from the LINEMOD dataset. The testing set encompasses 128 objects from Objaverse and 5 objects from LINEMOD. The images containing these objects are excluded from the training data, ensuring that all testing objects are previously unseen. All evaluated approaches are trained and tested on the same predefined image pairs, leading to a fair comparison.
@@
Table 4.2 provides the angular errors of the estimated relative object poses on the LINEMOD and Objaverse datasets. In the synthetic scenarios of Objaverse, DVMNet outperforms the previous methods by at least $7.92^{\circ}$ in terms of mean angular error. In the real scenarios of LINEMOD, DVMNet achieves the smallest angular error for most of the testing objects and reduces the mean angular error by at least $4.73^{\circ}$ compared to the other approaches. Moreover, we visualize the object pose depicted in the query image and show qualitative results in Fig. 7. The query object pose is determined as $\mathbf{R}_{q}=\Delta\mathbf{R}\mathbf{R}_{r}$, where $\Delta\mathbf{R}$ is the relative object pose and $\mathbf{R}_{r}$ denotes the object pose in the reference image. The ground-truth and predicted query object poses are represented as green and blue arrows, respectively. It is evident from Fig. 7 that the poses estimated by employing our DVMNet are consistently more similar to the ground truth than those obtained with the baselines.
@@
Figure 7: Qualitative results on LINEMOD [15]. We compute the object pose in the query image utilizing the relative object pose. The green and blue arrows represent the ground-truth object poses and the estimated ones, respectively.
@@ 4.4 Ablation Studies
Table 3: Effectiveness of the presented WCV algorithm. We report the mean angular errors on the CO3D dataset. The second row indicates the scenario where the WCV algorithm is replaced with a pose regression module. The third row presents a closest voxel algorithm without weights involved.
As a critical component of DVMNet, the presented WCV algorithm plays a pivotal role in achieving hypothesis-free and end-to-end relative object pose estimation. To substantiate the effectiveness of the WCV algorithm, we develop comprehensive ablation studies on the CO3D dataset.
@@
We first replace the WCV algorithm with a pose regression module. More concretely, we perform global average pooling over $\mathbf{V}_{q}$ and $\mathbf{V}_{r}$. The resulting feature embeddings are concatenated and passed through three fully connected layers to predict the 6D continuous representation of relative object rotation. We maintain all the other components in our framework unchanged to ensure a fair comparison. This alternative approach is also able to predict the relative object pose in a hypothesis-free and end-to-end fashion. However, as shown in Table 3, the mean angular error on the CO3D dataset increases by more than $10^{\circ}$ when the regression module is employed, showcasing the importance of the WCV algorithm in the presented hypothesis-free mechanism. Furthermore, we evaluate three counterparts of the WCV algorithm, i.e., a closest voxel algorithm without weights, a WCV algorithm with only replicated 2D object masks, and a WCV algorithm with only 3D objectness maps. The final weights of the voxel pairs are determined as $\mathbf{W}_{m}$ and $\mathbf{W}_{o}$ in the last two counterparts, respectively. The closest voxel algorithm delivers the worst results among these three variants, revealing that the pose estimation process is affected by the potential outliers. The optimal performance is achieved by leveraging both 2D object masks and 3D voxel objectness maps, which thus demonstrates the effectiveness of these components in DVMNet.
@@ 5 Conclusion
In this paper, we have introduced DVMNet, a novel deep voxel matching network for relative pose estimation of unseen objects. Unlike hypothesis-based approaches that rely on a multitude of discrete pose samples, DVMNet computes relative object poses in a hypothesis-free and end-to-end manner. We determine the object pose by solving a weighted least-squares problem based on voxel-voxel correspondences. The voxels are obtained by lifting the query and reference images to 3D via a dedicated autoencoder. We then align query voxels with the reference ones and apply a weighted closest voxel (WCV) algorithm for robust pose estimation. Comprehensive experiments on the CO3D, Objaverse, and LINEMOD datasets have demonstrated that our method excels in efficiently delivering accurate relative poses for unseen objects. In future work, we plan to explore the integration of DVMNet with zero-shot object detectors for 6D relative object pose estimation.
@@
Acknowledgment. This work was funded in part by the Swiss National Science Foundation via the Sinergia grant CRSII5-180359 and the Swiss Innovation Agency (Innosuisse) via the BRIDGE Discovery grant 40B2-0_194729.
[…]
Supplementary Material
6D Relative Object Pose Estimation
As introduced in the main paper, we are interested in extending the presented DVMNet to 6D relative object pose estimation in our future work. Notably, we followed the setting proposed in RelPose, RelPose++, and 3DAHV, assuming the availability of known object positions. Such information provides a strong prior, making the relation translation estimation less challenging. In this scenario, the primary challenge is the detection of an unseen object present in the query image. We achieve the zero-shot object detection by utilizing the detection module presented in Gen6D [21]. Specifically, the detection module predicts the 2D object center and 2D object scale, from which we compute the 3D object translation, following the implementation of Gen6D. The 3D object rotation is estimated by employing our DVMNet. It is worth noting that our goal is to achieve 6D object pose estimation in the single-reference setup, while Gen6D’s detection module utilizes dense-view references by default. Therefore, in our experiments, we feed a single reference to the detection module as the template.
We illustrate some preliminary results in Fig. 8 and Fig. 9, visualizing the 6D object pose in the query image as a 3D object bounding box. As shown in Fig. 8, the combination of Gen6D’s object detector and our DVMNet achieves promising results in some cases. However, in some scenarios shown in Fig. 9, Gen6D’s detector fails to effectively detect the unseen object depicted in the query image. Notably, the zero-shot object detection becomes more challenging in the single-reference setting due to the large pose difference between the query object and the reference. In this context, we plan to investigate the applicability of other zero-shot object detectors such as SAM [17] in future work.
Figure 8: Visualization of 6D pose estimation for unseen objects on LINEMOD [15]. The ground-truth 6D object pose and the predicted pose in the query image are depicted as green and blue 3D bounding boxes, respectively.
Figure 9: Failure cases of 6D pose estimation for unseen objects on LINEMOD [15].
Extension to Sparse-View References
  # References 1 2 3 4 5 6 7
SuperGlue 71.47 73.08 68.72 65.90 64.54 63.24 61.74
3DAHV 28.44 29.29 28.20 27.21 26.40 24.85 24.76
DVMNet 19.95 18.38 16.79 16.21 15.73 14.99 14.93
Table 4: Extension to sparse-view references. The experiment is conducted on CO3D [25] with the number of reference images varying from 1 to 7. The metric employed is the angular error between the computed query object pose and the ground truth.
As introduced in the literature [27, 47, 18], some downstream tasks such as 3D reconstruction often rely on sparse-view references. Intuitively, our DVMNet can be seamlessly integrated into these tasks, as it operates effectively with just a single reference. Therefore, we develop an experiment on the CO3D dataset, evaluating DVMNet with varying numbers of reference images, ranging from 1 to 7.
Specifically, given an unseen object during testing, we randomly sample $n$ images. These images are then fed into the presented DVMNet, with one image designated as the query and the remaining ones as references. The object pose in the query image is simply derived from the resulting $n-1$ relative object poses as
$\mathbf{R}_{q}=m(\frac{1}{n-1}\sum_{i=1}^{n-1}q(\Delta\mathbf{R}_{i}\mathbf{R}_{r}^{i})),$ (18)
where $\Delta\mathbf{R}_{i}$ and $\mathbf{R}_{r}^{i}$ denote the $i$-th relative object pose and reference pose, respectively, $q(\cdot)$ represents a function that converts a rotation matrix to the 6D continuous representation [53], and $m(\cdot)$ indicates the conversion from the 6D continuous representation to a rotation matrix. We also evaluate the representative image-matching (SuperGlue) and hypothesis-based (3DAHV) approaches in the sparse-view scenario. We ensure a fair comparison by utilizing the same strategy of query object pose estimation for these methods.
We report the resulting angular errors in Table 4. It is evident that (i) the angular error of our DVMNet decreases as more reference images are involved, and (ii) DVMNet consistently yields the smallest angular error. This observation demonstrates the promising compatibility of our approach with sparse-view reference images.
Robustness to Occlusions
  Method SuperGlue 3DAHV DVMNet
Angular Error $\downarrow$ 73.72 51.49 49.02
Table 5: Experimental results on the LINEMOD-O [5] dataset. The mean angular errors are reported.
Given that object pose estimation is often challenged by occlusions, we assess the robustness in scenarios involving occlusions by conducting an experiment on the LINEMOD-O [5] dataset. The testing data comprises three unseen objects, i.e., cat, driller, and duck. We report the mean angular errors of the evaluated methods in Table 5. Our DVMNet outperforms both the image-matching method, SuperGlue, and the hypothesis-based method, 3DAHV, showcasing better robustness against occlusions.
  Method Acc @ $30^{\circ}$ (%) $\uparrow$
RelPose 64.2
RelPose++ 77.0
3DAHV 83.5
PoseDiffusion 81.8
DVMNet 84.7
Table 6: Additional experimental results on CO3D. Acc @ $30^{\circ}$ is employed as a metric.
Additional Results on CO3D
As listed in Tabel 6, we report more results on CO3D, using Acc @ $30^{\circ}$ as a metric. Note that PoseDiffusion [37] takes multiple views ($>2$) as input by default, while all the other evaluated methods employ two views. For a fair comparison, we evaluate PoseDiffusion on CO3D using two views. Moreover, The pose parameters are iteratively updated 100 times during the denoising process in PoseDiffusion, making the method time-consuming compared with our single forward pass mechanism.

### v2

@@ 4.3 Relative Object Rotation Estimation on GROP
Recently, a new benchmark called GROP for relative rotation estimation of unseen objects was introduced in [14]. This benchmark comprises two datasets, i.e., Objaverse [24] and LINEMOD [18]. Both synthetic and real images with diverse object poses are considered. We perform experiments on these two datasets, following the same setup as described in [14]. More concretely, the synthetic images are generated by rendering the object models of the Objaverse dataset from different viewpoints [62]. Several sequences of calibrated real images that depict 13 texture-less household objects are provided from the LINEMOD dataset. The testing set encompasses 128 objects from Objaverse and 5 objects from LINEMOD. The images containing these objects are excluded from the training data, ensuring that all testing objects are previously unseen. All evaluated approaches are trained and tested on the same predefined image pairs, leading to a fair comparison. As in [14], we crop the object from the query image employing the ground-truth object bounding boxes.
@@
Table III provides the angular errors of the estimated relative object rotations on the LINEMOD and Objaverse datasets. In the synthetic scenarios of Objaverse, DVMNet++ outperforms the previous methods by at least $7.92^{\circ}$ in terms of mean angular error. In the real scenarios of LINEMOD, DVMNet++ achieves the smallest angular error for most of the testing objects and reduces the mean angular error by at least $4.73^{\circ}$ compared to the other approaches. Moreover, we visualize the object rotation depicted in the query image and show qualitative results in Fig. 9. The query object rotation is determined as $\mathbf{R}_{q}=\Delta\mathbf{R}\mathbf{R}_{r}$. The ground-truth and predicted query object rotations are represented as green and blue arrows, respectively. It is evident from Fig. 9 that the rotations estimated with our DVMNet++ are consistently more similar to the ground truth than those obtained with the baselines.
@@
Figure 9: Qualitative rotation estimation results on LINEMOD [18]. We visualize the object rotation in the query image based on the relative object rotation. The green and blue arrows represent the ground-truth object rotations and the estimated ones, respectively.
4.4 Relative Object Pose with Open-Set Detector
Recall that, in the preceding experiments, we use the ground-truth object bounding box to crop the object from the query image, which aligns with the setting in [12, 13, 14]. To evaluate the effectiveness of the proposed open-set object detector, we conduct experiments on CO3D and LINEMOD, replacing the ground-truth bounding boxes with the predicted ones. We use the object category as the text prompt when performing the object detection on CO3D. Since LINEMOD contains cluttered scenarios, unseen object detection is more challenging. Therefore, we incorporate additional attribute descriptions, such as color and material, based on the reference image.
We evaluate our method in terms of translation estimation error (Trans. Error) and rotation estimation error (Rota. Error). Specifically, we obtain the object bounding box parameters using the presented object detector. The relative object translation is computed based on Eq. 1. The relative object rotation is estimated using the cropped query image and the reference image. To eliminate the scale ambiguity, we compute the angular error [38] of the translation, which is defined as
$err_{T}=\frac{\Delta\hat{\mathbf{T}}\cdot\Delta\mathbf{T}_{gt}}{{||\Delta\hat{\mathbf{T}}||}_{2}{||\Delta\mathbf{T}_{gt}||}_{2}},$ (21)
where $\Delta\hat{\mathbf{T}}$ and $\Delta\mathbf{T}_{gt}$ denote the predicted relative object translation and the ground truth, respectively. The experimental results on CO3D and LINEMOD are reported in Table IV and Table V, respectively. Our open-set object detector yields highly accurate relative object translation estimates, with a mean translation error of $6.47^{\circ}$ on CO3D and $1.55^{\circ}$ on LINEMOD. This demonstrates that, on average, the predicted bounding box center is close to the ground-truth object center in the query image. Moreover, the rotation error based on the predicted object bounding boxes is comparable to those obtained with the ground-truth object positions. This observation highlights the effectiveness of our approach in two aspects: First, our deep voxel matching network is robust to noise in object detection; second, our open-set object detector provides reliable object position parameters. We showcase some detection results in Fig. 10. As shown in this figure, the object proposals generated based on the natural language understanding are noisy, particularly in cluttered scenarios. Our detector building upon multiple modalities is capable of identifying the correct bounding box from the candidates.
We also evaluate the efficiency of the open-set object detector. On average, the detection takes $\emph{92}ms$ on the CO3D dataset. This demonstrates that our method delivers reliable object detection results for relative object pose estimation with acceptable computational overhead.
TABLE IV: Unseen object detection results on CO3D [23]. We replace the ground-truth object bounding boxes with the ones predicted by our open-set object detector. The translation estimation errors (Tran. Error) and rotation estimation errors (Rota. Error) with the predicted object bounding boxes are reported.
  CO3D Ball Book Couch Frisb. Hotd. Kite Remot. Sandw. Skate. Suitc. Mean
Tran. Error 4.23 3.44 3.05 10.33 2.58 3.13 4.62 2.63 2.63 3.38 6.47
Rota. Error 27.66 22.04 19.44 22.69 19.98 16.67 11.47 18.91 20.25 17.25 19.64
TABLE V: Unseen object detection results on LINEMOD [18]. We maintain the same experimental settings as those used on CO3D. We show the translation and rotation estimation errors, using the proposed object detector in our pipeline.
  LINEMOD Cat Ben. Cam. Dri. Duck Mean
Tran. Error 1.04 1.93 1.44 1.64 1.67 1.55
Rota. Error 33.62 38.56 35.22 48.31 43.24 39.79
Figure 10: Qualitative unseen object detection results on CO3D [23] and LINEMOD [18]. The object proposals are obtained by employing the open-vocabulary object detector [21]. We determine the final detection result from the proposals using the introduced image retrieval strategy.
@@ 4.5 Ablation Studies
4.5.1 Weighted Closest Voxel Algorithm
TABLE VI: Effectiveness of our WCV algorithm. We report the mean angular errors of relative object rotation estimation on the CO3D [23] dataset. The second row indicates the scenario where the WCV algorithm is replaced with a rotation regression module. The third row presents the closest voxel algorithm without weights involved.
@@ 4.5.1 Weighted Closest Voxel Algorithm
As a critical component of our deep voxel matching network, the weighted closest voxel (WCV) algorithm plays a pivotal role in achieving hypothesis-free and end-to-end relative object rotation estimation. To substantiate the effectiveness of the WCV algorithm, we develop comprehensive ablation studies on the CO3D dataset. We first replace the WCV algorithm with a rotation regression module. More concretely, we perform global average pooling over $\mathbf{V}_{q}$ and $\mathbf{V}_{r}$. The resulting feature embeddings are concatenated and passed through three fully connected layers to predict the 6D continuous representation of relative object rotation. We maintain all the other components in our framework unchanged to ensure a fair comparison. This alternative approach is also able to predict the relative object rotation in a hypothesis-free and end-to-end fashion. However, as shown in Table VI, the mean angular error on the CO3D dataset increases by more than $10^{\circ}$ when the regression module is employed, showcasing the importance of the WCV algorithm in the presented hypothesis-free mechanism. Furthermore, we evaluate three counterparts of the WCV algorithm, i.e., a closest voxel algorithm without weights, a WCV algorithm with only replicated 2D object masks, and a WCV algorithm with only 3D objectness maps. The final weights of the voxel pairs are determined as $\mathbf{W}_{m}$ and $\mathbf{W}_{o}$ in the last two counterparts, respectively. The closest voxel algorithm delivers the worst results among these three variants, revealing that the rotation estimation process is affected by the potential outliers. The best performance is achieved by leveraging both 2D object masks and 3D voxel objectness maps, which thus demonstrates the effectiveness of these components in the proposed rotation estimation module.
TABLE VII: Extension to sparse-view references. The experiment is conducted on CO3D [23] with the number of reference images varying from 1 to 7. We report the angular error between the computed query object rotation and the ground truth.
  # References 1 2 3 4 5 6 7
SuperGlue [16] 71.47 73.08 68.72 65.90 64.54 63.24 61.74
3DAHV [14] 28.44 29.29 28.20 27.21 26.40 24.85 24.76
DVMNet++ 19.95 18.38 16.79 16.21 15.73 14.99 14.93
4.5.2 Extension to Sparse References
Note that, by default, we assume that a single reference image is available in our experiments. However, to account for scenarios where multiple reference images may be provided, we conduct an experiment on the CO3D dataset to evaluate the compatibility of our method with such sparse references. Specifically, given an unseen object during testing, we randomly sample $n$ images, with $n$ ranging from 2 to 8. These images are then fed into our network, with one image designated as the query and the remaining ones as references. As shown in the preceding experiments, relative object rotation estimation is more challenging than translation estimation. Consequently, we focus on evaluating rotation estimation in this experiment. The object rotation in the query image is simply derived from the resulting $n-1$ relative object rotations as
$\mathbf{R}_{q}=h^{-1}(\frac{1}{n-1}\sum_{i=1}^{n-1}h(\Delta{\hat{\mathbf{R}}_{i}}\mathbf{R}_{r}^{i})),$ (22)
where $\Delta{\hat{\mathbf{R}}_{i}}$ and $\mathbf{R}_{r}^{i}$ denote the $i$-th relative object rotation and reference rotation, respectively, $h(\cdot)$ represents a function that converts a rotation matrix to the 6D continuous representation [37], and $h^{-1}(\cdot)$ indicates the inverse conversion. We also evaluate the representative image-matching (SuperGlue) and hypothesis-based (3DAHV) approaches in the sparse-view scenario. We ensure a fair comparison by utilizing the same strategy of query object rotation estimation for these methods. We report the resulting rotation errors in Table VII. It is evident that (i) the rotation error of our method decreases as more reference images are involved, and (ii) our method consistently yields the smallest rotation error. These observations demonstrate the promising compatibility of our approach with sparse reference images.
TABLE VIII: Effectiveness of the object detector. The relative object translation is approximated using the parameters of the bounding box. GT indicates that the ground-truth bounding box is used. w/o RGB means the detection result is selected from the proposals using the confidence scores. We report the mean translation errors on LINEMOD [18].
  Method GT w/o RGB Ours
Tran. Error 0.75 5.27 1.55
4.5.3 Open-Set Object Detection
To shed more light on the effectiveness of our open-set object detection module, we conduct experiments on LINEMOD, comparing the method with several alternatives. We first evaluate the method (GT) using the ground-truth object bounding box. Recall that we approximate the 2D projection of the 3D object center in the object coordinate system as the center of the object bounding box. These positions may differ, even when using the ground-truth bounding box. In this context, the translation error of GT reflects the systematic error introduced by the approximation. Moreover, we remove the image-retrieval module from the detection framework. We utilize the confidence scores predicted by the open-vocabulary detector [21] to identify the detection result from the proposals. As listed in Table VIII, this alternative (w/o RGB) leads to more erroneous translation estimations, which demonstrates the effectiveness of our multi-modal object detector.
TABLE IX: Experimental results on the LINEMOD-O [25] dataset. We report the mean angular errors.
  Method SuperGlue 3DAHV DVMNet++
Tran. Error - - 6.47
Rota. Error 73.72 51.49 48.82
4.5.4 Robustness to Occlusions
Given that object pose estimation is often challenged by occlusions, we assess robustness in scenarios involving occlusions by conducting an experiment on the LINEMOD-O [25] dataset. The testing data comprises three unseen objects, i.e., cat, driller, and duck. We report the mean angular errors in Table IX. The results showcase the promising robustness of our method to occlusions.
@@ 5 Conclusion
In this paper, we have introduced DVMNet++, a novel approach for relative pose estimation of unseen objects. Given a single RGB image as the reference, DVMNet++ identifies the object in the query image and computes the relative object pose without relying on the GT object bounding box or rotation hypotheses. This has been achieved via a multi-modal open-set object detector and a deep voxel matching network. Comprehensive experiments on the CO3D, Objaverse, LINEMOD, and LINEMOD-O datasets have demonstrated that our DVMNet++ excels in efficiently delivering accurate relative poses for previously unseen objects.
@@
Acknowledgments. This work was funded in part by the Swiss National Science Foundation via the Sinergia grant CRSII5-180359 and the Swiss Innovation Agency (Innosuisse) via the BRIDGE Discovery grant 40B2-0_194729.